Giac collected money in his piggy bank until he was 10 years old. On his birthday, he deposited all $213.45 in the bank. His account earned 9% interest, compounded quarterly. On his 18th birthday, Giac took all of his money out to help buy a car. How much money was in the account at that time?
This is an applied problem using the rate of growth.

Answer by Theo(13342)

on his 10th birthday, he deposited 213.45 in the bank.
his account earned 9% per year compounded quarterly.
on his 18th birthday, he took the money out to help buy a car.

his money was in the bank for 8 years.

9% per year is equal to 9/100 per year = .09 per year.

time periods are in months because the interest rate is compounded monthly, therefore:

interest rate per month = .09/12 = .0075 per month.
number of months = 8 * 12 = 96 months.

formula is f = p * (1+r)^n

f is the future value
p is the present value = 213.45
r is the interest rate per month = .0075 per month.
n is the number of months = 96 months.

formula becomes:

f = 213.45 * (1.0075)^96

solve for f to get f = 437.3422362.

round to 2 decimal places to get f = $437.34.

you can use your calculator to confirm this is true, based on the future value of a present amount formula shown above.

there are other financial formulas you might find useful along with a reference to a financial calculator that should be helpful to confirm you did the problems correctly.

they're contained in the following reference.

https://www.algebra.com/algebra/homework/Finance/THEO-2016-04-29.lesson#notes